OpenStudy (anonymous):

Solve for y in the equation –11y = –143. A. –13 B. –12 C. 12 D. 13

OpenStudy (anonymous):

@Astrophysics

OpenStudy (agreene):

remember that to isolate a variable from a constant you need to divide it off from both sides.

OpenStudy (anonymous):

Hm...?

OpenStudy (agreene):

\[ax=b\] \[\frac{ax}a=\frac{b}a\] \[x=\frac{b}a\]

OpenStudy (astrophysics):

Hey so for this problem, you want to isolate y, meaning get y on its own. To do this you have to get rid of the -11, since we are multiplying -11 here by y we have to do it's opposite which is division! And that will cancel it out. \[\frac{ -11y }{ -11 } = \frac{ -143 }{ -11 }\] You see I did it to both sides, because what you do to one side, you must also do it to the other. So we can cancel the -11/-11 on the left because that just = 1 and 1*y = y, and now we just divide as you would normally on the right side to get 13, remember negative divided by negative = positive and negative multiplied by negative is also positive. \[y = 13\]

OpenStudy (anonymous):

Oh I get it now!

OpenStudy (astrophysics):

Great! :)